OpenStudy (anonymous):

How do you solve: |x-2| = 5? I will post answer choices.

OpenStudy (anonymous):

x = 7..

OpenStudy (mertsj):

|x-2|=5 means that the numbers are exactly 5 untis from 0.

OpenStudy (anonymous):

A) {-3,-7} B) {-7} C) {-3,7} D) No solution

OpenStudy (anonymous):

solve two equation, plus or minus

OpenStudy (anonymous):

But it doesnt have 7 in any of them.

OpenStudy (anonymous):

the answer is -7

OpenStudy (mertsj):

The numbers that are exactly 5 units from 0 are 5 and -5 so write the two equations: x-2=5 or x-2 = -5

OpenStudy (mertsj):

Solve each equation.

OpenStudy (anonymous):

So I got confused :/ How is it -7? Cuz it is 5+2=7.

OpenStudy (mertsj):